[6-a 4] * [-2] = [2a]
[-3 -3(d-2] * [3] = [-3]

find a and d... this is a matrice

Answer by stanbon(75887)

[6-a 4] * [-2] = [2a]
[-3 -3(d-2] * [3] = [-3]
----------------------------------------
(6-a)(-2) + 4*3 = 2a
-12+2a + 12 = 2a
2a = 2a
a is indeterminate
-------------------------
-3*-2 + -2(d-2)(3) = -3
6 - 6(d-2) = -3
-6d+12 = -9
-6d = -21
d = 7/2
-------------------------
Cheers,
Stan H.